# ATOM: AdapTive and OptiMized dynamic temporal knowledge graph construction using LLMs

### v1

@@ 4.4 TKG stability
Rerunning the same prompt multiple times leads to variations in output. To evaluate the stability score, 5-tuples are initially extracted from both atomic facts (denoted as (F) in Table 1) and lead paragraphs individually, establishing a baseline named Run 1. This extraction process is subsequently repeated twice more without altering any parameters. As shown in Table 1, centroids of the extracted 5-tuples’ embeddings from atomic facts remain nearly constant across runs 1 and 2, exhibiting very low standard deviations and high stability scores. In contrast, the centroids of the extracted 5-tuples’ embeddings from the lead paragraphs demonstrate greater variability. This addresses (C2) and reflects the effect of atomic facts in maintaining a stable construction of TKGs with a gain of $\sim 17\%$.
Table 1: Stability $S_{r}$ evaluated by rerunning the 5-tuples extraction process multiple times without any modifications, using gpt-4.1-2025-04-14 with Run 1 as a baseline. The extraction is performed on (F) atomic facts and (L) lead paragraphs.
@@ 4.5 The exhaustivity and quality of the 5-tuples
5-tuples are extracted from both atomic facts (denoted as (F) in Table 2) and lead paragraphs (L). As shown in Table 2, higher values are maintained by the $R_{\text{MATCH}}$ and $R_{\text{MATCH}_{t}}$ rates when atomic facts are provided as input compared to when lead paragraphs are used, with a gain of $\sim 18\%$ achieved on temporal exhaustivity and $\sim 31\%$ on factual exhaustivity, along with a gain of approximately $\sim 31\%$ on factual omission. Hence, atomic fact decomposition improves temporal and factual exhaustivity, and this addresses (C1). However, an increase in $R_{\text{HALL}}$ by $\sim 9\%$ is observed, which is attributed to the addition of ’inferred’ atomic facts from the lead paragraphs by the LLM during the atomic fact decomposition. Consequently, the LLM extracts 5-tuples that do not exist in the gold standard. Furthermore, a higher $R_{\text{OM}_{t}}$ is observed, which is attributed to imperfections that may occur during atomic fact decomposition, where temporal information may not be assigned to atomic facts by the LLM. This error propagates to 5-tuples extraction in $R_{\text{OM}_{t}}$. It is a trade-off that is discussed further in Section 6.
Table 2: 5-tuple quality metrics. The extraction is performed on (F) atomic facts and (L) lead paragraphs.
@@
$R_{\text{MATCH}}$ 0.405 $\pm$ 0.150 0.720 $\pm$ 0.143
   $R_{\text{MATCH}_{t}}$ 0.176 $\pm$ 0.123 0.354 $\pm$ 0.165
   $R_{\text{OM}_{t}}$ 0.229 $\pm$ 0.131 0.366 $\pm$ 0.158
   $R_{\text{HALL}_{t}}$ 0.000 $\pm$ 0.000 0.000 $\pm$ 0.000
$R_{\text{OM}}$ 0.595 $\pm$ 0.150 0.280 $\pm$ 0.143
$R_{\text{HALL}}$ 0.333 $\pm$ 0.172 0.428 $\pm$ 0.128
@@ 4.6 ATOM’s time complexity
Given the demonstrated benefits of atomic fact decomposition in improving exhaustivity and stability, all subsequent experiments utilize atomic facts as input rather than lead paragraphs. All baseline methods are run using gpt-4.1-2025-04-14. ATOM employs 8 threads and a batch size of 40 atomic facts for 5-tuples extraction, which respects OpenAI rate limits. iText2KG and Graphiti separate entity and relation extraction, increasing latency. Graphiti’s incremental entity/relation resolution, which relies on the LLM, limits parallel requests and significantly increases latency as the graph expands. Similarly, iText2KG is incremental, restricting parallel requests. Although iText2KG uses a distance metric for resolution, reducing LLM dependency, its separate extraction steps double the number of LLM calls and induce isolated entities that require further LLM iterations. Conversely, ATOM’s architecture facilitates (1) parallel LLM calls, (2) parallel merge of atomic TKGs, (3) LLM-independent merging, and (4) temporal resolution. This design reduces latency by 93.8% compared to Graphiti and 95.3% compared to iText2KG (Figure 3). ATOM’s Module-3 accounts for only 13% of its total latency, with the remainder attributed to API calls, which can be further minimized through either increasing the batch size (by upgrading the API tier) or scaling hardware for local LLM deployment.
@@
4.7 ATOM’s DTKG construction
Table 3 shows that ATOM and iText2KG demonstrate comparable entity and relation resolution performance, as both employ a distance metric for merging. ATOM shows an improvement over Graphiti, whose incremental, LLM-based entity and relation resolution degrades with graph expansion (increasing context size), which is consistent with findings in Section 4.3. Beyond entity and relation resolution, temporal resolution reveals more significant differences between ATOM and Graphiti. The examples in Figures F.3 and F.4 in the Appendices illustrate atomic facts observed at different times that refer to the same information but with different validity periods. In these examples, Graphiti creates separate relations for 5-tuples with different validity periods ($t_{start}$, $t_{end}$), while ATOM detects these similar relations and extends their validity period history. This temporal resolution serves two functions: tracking events that naturally appear and disappear over time, and matching relations with only $t_{start}$ or only $t_{end}$ to complete their validity periods as additional information becomes available. Additionally, Graphiti incorporates validity periods only and does not allow for observation time modeling, treating observation time as $t_{start}$. This can induce errors. For example, if a news article observed on "January 23, 2020" states "The mysterious respiratory virus spread to at least 10 other countries," Graphiti would set $t_{start}=23\text{-}01\text{-}2020$, while the statement does not specify a validity period and the true validity could be weeks or days before its publication. In contrast, ATOM separately models observation time and validity periods, allowing it to recognize atomic facts without a validity period and avoid incorrect temporal assignments (Example in Figure F.2 in the Appendices).
@@ 4.7 ATOM’s DTKG construction
Precision-ER 0.994 0.967 0.974
Recall-ER 0.993 0.952 0.980
F1-Score-ER 0.994 0.959 0.977
Precision-$R$ 1 0.917 0.991
Recall-$R$ 1 0.888 0.988
F1-Score-$R$ 1 0.902 0.989
@@ 6 Limitations
ATOM has some limitations that warrant consideration. First, the atomic fact decomposition can introduce error propagation: the LLM may generate inferred facts not present in the source text, leading to increased hallucination rates, and may fail to properly assign temporal information to atomic facts, resulting in temporal omissions in the extracted 5-tuples (Section 4.5). A potential improvement consists of fine-tuning an LLM model specifically for this task. Second, the distance-metric-based merging approach, while scalable and efficient, can occasionally merge semantically distinct named entities that exhibit high similarity (e.g., gpt-5:model and gpt-3.5:model). A supervised entity/relation resolution classifier trained on labeled entity pairs could replace the threshold-based approach.
@@ Appendix B Example of the atomic fact decomposition
  • Real Madrid won the Champions League final match on June 18, 2024. (Observation $t_{obs}=[01-01-2025]$, $t_{start}=[18-06-2024]$, $t_{end}=[.]$)
  • The Champions League final match ended with a 2-1 victory for Real Madrid on June 18, 2024. (Observation $t_{obs}=[01-01-2025]$, $t_{start}=[18-06-2024]$, $t_{end}=[.]$)
  • Fans of Real Madrid celebrated the Champions League final match victory across the city on June 18, 2024. (Observation $t_{obs}=[01-01-2025]$, $t_{start}=[18-06-2024]$, $t_{end}=[.]$)
@@ Appendix C Estimating the merging thresholds
Figure F.1: Example overview of ATOM’s pipeline. It begins with atomic fact decomposition, followed by the extraction of atomic TKGs from these facts, which are then merged in parallel. When an incoming update arrives, ATOM handles the temporal resolution by transforming the end action into the affirmative part while modifying only the $t_{end}$, then merges the resulting atomic TKG with the existing DTKG.
@@ Appendix D Observation time modeling
Figure F.2: Two DTKGs constructed using ATOM and Graphiti from 09-01-2020 (in UNIX, 1578524400) to 23-01-2020 (in UNIX, 1579734000) from 2020-COVID-NYT dataset. Left (ATOM): Preserves observation times ($t_{obs}$) separately from validity periods, with timestamps encoded in UNIX format to eliminate overhead associated with string parsing operations and timezone conversion calculations. Right (Graphiti): Treats observation time as validity start time. valid_atvalid\_at corresponds to $t_{start}$ in Graphiti’s time modeling. The highlighted fact “The mysterious respiratory virus spread to at least 10 other countries” is observed on 23-01-2020, but this does not guarantee the spread occurred at that time. ATOM’s dual-time modeling prevents such temporal misattribution.
@@
Figure F.3: Temporal resolution comparison between ATOM and Graphiti. Two atomic facts observed on January 28, 2020, report death counts from January 24 (26 deaths) and January 27 (at least 80 deaths). Left (ATOM): performs temporal resolution by detecting similar relations and extending their validity period history ($t_{end}$ in the figure). Right (Graphiti): creates separate relations for each atomic fact, resulting in duplication. Moreover, Graphiti misinterprets “By January 24, 2020” and “By January 27, 2020” as validity start times rather than validity end times, leading to temporal misattribution.
@@
Figure F.4: Temporal resolution comparison between ATOM and Graphiti. Two atomic facts observed on different dates (April 16 and April 19, 2020) describe protest activities during two time periods (“the week of April 13” and “the week of April 19”). Left (ATOM): merges similar relations and extend their validity periods ($t_{start}$ and $t_{end}$ in the figure). Right (Graphiti): maintains separate relations for each atomic fact. Moreover, Graphiti fails to translate “In the week of April 13, 2020” and “In the week of April 19, 2020” into proper validity periods as ATOM does.

### v2

@@ 4.4 TKG stability
Rerunning the same prompt multiple times leads to variations in output. To evaluate the stability score, 5-tuples are initially extracted from both atomic facts (denoted as (F) in Table 1) and lead paragraphs individually, establishing a baseline named Run 1. This extraction process is subsequently repeated twice more without altering any parameters. As shown in Table 1, the 5-tuple sets extracted from atomic facts exhibit high Jaccard similarity with the baseline across runs. In contrast, the extraction from lead paragraphs demonstrates significantly lower overlap, revealing high variability. This addresses (C2) and reflects the effect of atomic facts in maintaining a stable construction of TKGs with a gain of $\sim 33\%$.
Table 1: Stability $S_{r}$ evaluated by rerunning the 5-tuple extraction process multiple times without any modifications, using gpt-4.1-2025-04-14 with Run 1 as a baseline. The extraction is performed on (F) atomic facts and (L) lead paragraphs.
@@ 4.5 The exhaustivity and quality of the 5-tuples
5-tuples are extracted from both atomic facts (denoted as (F) in Table 2) and lead paragraphs (L). As shown in Table 2, ATOM significantly outperforms the baseline in exhaustivity, with factual recall $R_{f}$ increasing by $\sim 31\%$ and factual-temporal recall $R_{f,t}$ by $\sim 18\%$. These gains lead to a substantial improvement in overall extraction quality, resulting in increased factual $F_{1,f}$ and factual-temporal $F_{1,f,t}$ scores. Hence, atomic fact decomposition effectively mitigates information loss, addressing (C1). However, this improved exhaustivity presents a trade-off: a decrease in factual precision $P_{f}$ of $\sim 9\%$. This occurs because the baseline (L) acts conservatively, extracting only the most obvious facts, whereas ATOM’s decomposition forces the LLM to explicitize context. This process can generate ’inferred’ atomic facts that are semantically plausible but not strictly present in the gold standard, leading to false positives. Furthermore, a marginal decrease of $\sim 1\%$ in factual-temporal precision $P_{f,t}$ is observed. This is attributed to imperfections during the decomposition phase, where the LLM may fail to assign temporal information to certain atomic facts. Consequently, this error propagates to the 5-tuple extraction. This trade-off is discussed further in Section 6.
Table 2: 5-tuple quality metrics. The extraction is performed on (L) lead paragraphs and (F) atomic facts.
@@
$P_{f}$ 0.667 $\pm$ 0.172 0.572 $\pm$ 0.128
$R_{f}$ 0.405 $\pm$ 0.150 0.720 $\pm$ 0.143
$F_{1,f}$-score 0.504 0.638
$P_{f,t}$ 0.292 $\pm$ 0.182 0.281 $\pm$ 0.130
$R_{f,t}$ 0.176 $\pm$ 0.123 0.354 $\pm$ 0.165
$F_{1,f,t}$-score 0.220 0.313
@@ 4.6 ATOM’s time complexity
Given the demonstrated benefits of atomic fact decomposition in improving exhaustivity and stability, all subsequent experiments utilize atomic facts as input rather than lead paragraphs. All baseline methods are run using gpt-4.1-2025-04-14. ATOM employs 8 threads and a batch size of 40 atomic facts for 5-tuple extraction, which respects OpenAI rate limits. iText2KG and Graphiti separate entity and relation extraction, increasing latency. Graphiti’s incremental entity/relation resolution, which relies on the LLM, limits parallel requests and significantly increases latency as the graph expands. Similarly, iText2KG is incremental, restricting parallel requests. Although iText2KG uses a distance metric for resolution, reducing LLM dependency, its separate extraction steps double the number of LLM calls and induce isolated entities that require further LLM iterations. Conversely, ATOM’s architecture facilitates (1) parallel LLM calls, (2) parallel merge of atomic TKGs, (3) LLM-independent merging, and (4) temporal resolution. This design reduces latency by 93.8% compared to Graphiti and 95.3% compared to iText2KG (Figure 4). ATOM’s Module-3 accounts for only 13% of its total latency, with the remainder attributed to API calls, which can be further minimized through either increasing the batch size (by upgrading the API tier) or scaling hardware for local LLM deployment.
@@
4.7 Consistency of ATOM’s DTKG construction
Table 3 shows that ATOM and iText2KG demonstrate comparable entity and relation resolution performance, as both employ a distance metric for merging. ATOM shows an improvement over Graphiti, whose incremental, LLM-based entity and relation resolution degrades with graph expansion (increasing context size), which is consistent with findings in Section 4.3. Beyond entity and relation resolution, temporal resolution reveals more significant differences between ATOM and Graphiti. The examples in Figures F.2 and F.3 in the Appendices illustrate atomic facts observed at different times that refer to the same information but with different validity periods. In these examples, Graphiti creates separate relations for 5-tuples with different validity periods ($t_{\text{start}}$, $t_{\text{end}}$), while ATOM detects these similar relations and extends their validity period history. This temporal resolution serves two functions: tracking events that naturally appear and disappear over time, and matching relations with only $t_{\text{start}}$ or only $t_{\text{end}}$ to complete their validity periods as additional information becomes available. Additionally, Graphiti incorporates validity periods only and does not allow for observation time modeling, treating observation time as $t_{\text{start}}$. This can induce errors. For example, if a news article observed on "January 23, 2020" states "The mysterious respiratory virus spread to at least 10 other countries," Graphiti would set $t_{\text{start}}=23\text{-}01\text{-}2020$, while the statement does not specify a validity period and the true validity could be weeks or days before its publication. In contrast, ATOM separately models observation time and validity periods, allowing it to recognize atomic facts without a validity period and avoid incorrect temporal assignments (Example in Figure F.1 in the Appendices).
@@ 4.7 Consistency of ATOM’s DTKG construction
$P_{ER}$ 0.994 0.967 0.974
$R_{ER}$ 0.993 0.952 0.980
$F_{1,ER}$-score 0.994 0.959 0.977
$P_{RR}$ 1 0.917 0.991
$R_{RR}$ 1 0.888 0.988
$F_{1,RR}$-score 1 0.902 0.989
@@ 6 Limitations
ATOM has some limitations that warrant consideration. First, the atomic fact decomposition can introduce error propagation: the LLM may generate inferred facts not present in the source text, leading to increased hallucination, and may fail to properly assign temporal information to atomic facts, resulting in a decrease in the factual-temporal precision (Section 4.5). A potential improvement consists of fine-tuning an LLM model specifically for this task. Second, the distance-metric-based merging approach, while scalable and efficient, can occasionally merge semantically distinct named entities that exhibit high similarity (e.g., gpt-5:model and gpt-3.5:model). A supervised entity/relation resolution classifier trained on labeled entity pairs could replace the threshold-based approach. Third, ATOM entails higher costs than direct extraction (Table T.2 in the Appendices) due to the added decomposition step and the increased output tokens required to maintain high exhaustivity. This cost could be mitigated by using or fine-tuning open-weight models specifically for extracting 5-tuples from atomic facts.
7 Ethics statement
We utilized AI assistants exclusively for grammatical corrections and text polishing. All scientific claims, experimental designs, and original ideas are solely those of the authors.
@@ Appendix B Example of the atomic fact decomposition
  • Real Madrid won the Champions League final match on June 18, 2024. (Observation $t_{obs}=[01-01-2025]$, $t_{\text{start}}=[18-06-2024]$, $t_{\text{end}}=[.]$)
  • The Champions League final match ended with a 2-1 victory for Real Madrid on June 18, 2024. (Observation $t_{obs}=[01-01-2025]$, $t_{\text{start}}=[18-06-2024]$, $t_{\text{end}}=[.]$)
  • Fans of Real Madrid celebrated the Champions League final match victory across the city on June 18, 2024. (Observation $t_{obs}=[01-01-2025]$, $t_{\text{start}}=[18-06-2024]$, $t_{\text{end}}=[.]$)
@@ Appendix D Observation time modeling
Appendix E Annotation guidelines for 2020-COVID-NYT
The ground truth for the 2020-COVID-NYT dataset was constructed through a two-step semi-manual annotation process utilizing the Claude Sonnet 4 chat model with extended thinking [^11] https://claude.ai/. We selected this model after observing its superior performance on a small subset of lead paragraphs compared to other available SOTA LLMs. Human annotators supervised the process, provided with the lead paragraph and an associated observation date ($t_{obs}$), and were instructed to follow the guidelines detailed below.
E.1 Phase 1: atomic fact decomposition
In this phase, the model decomposed complex paragraphs into atomic, self-contained, and temporally grounded statements. Annotators were tasked with verifying and correcting the exhaustivity of the decomposition, ensuring the absence of hallucinatory facts, and validating the correctness of validity times for temporal facts, strictly following these guidelines:
Atomicity:
Compound sentences must be split. Each atomic fact must contain exactly one piece of information. Redundancies and duplicated information were removed.
Decontextualization:
Facts must be understandable in isolation. Pronouns (e.g., "he", "it") were replaced with full entity names, and necessary modifiers were included.
Temporal normalization:
All relative time references were converted to absolute dates based on the $t_{obs}$. For example: Today $\rightarrow$ exact $t_{obs}$;  Yesterday $\rightarrow$ $t_{obs}-1$ day;  Last week $\rightarrow$ Monday of the week preceding $t_{obs}$;  This year $\rightarrow$ January 1st of the $t_{obs}$ year.
End actions:
If the text indicated the end of a role or action (e.g., "leaving a position"), the termination was captured along with its specific timestamp.
E.2 Phase 2: 5-tuple extraction
In the second phase, the model extracted temporal 5-tuples $(e_{s},r,e_{o},t_{\text{start}},t_{\text{end}})$ from the atomic facts generated in Phase 1. Similarly, annotators verified and corrected the extracted tuples for exhaustivity, factual accuracy, and strict temporal validity, ensuring adherence to the following guidelines:
Entity Annotation
Entities were defined as distinct concepts (e.g., Person, Organization, Position). Dates were strictly excluded from entity mentions (they are reserved for temporal arguments).
Relationship and temporal annotation
Relations were extracted to capture the interaction between entities, with strict temporal rules:
  • Canonical present tense: All relation names were annotated in the present tense (e.g., is_CEO, works_at) regardless of whether the event was past, present, or future. This enhances semantic consistency across the TKG.
  • Validity periods: Temporal information was mapped strictly to $t_{\text{start}}$ (validity start) and $t_{\text{end}}$ (validity end) based on the absolute dates resolved in Phase 1.
    • Affirmative actions (e.g., "became CEO") populated $t_{\text{start}}$.
    • End actions (e.g., "is no longer the CEO") populated $t_{\text{end}}$, while the relation name remained in the affirmative (e.g., is_CEO).
@@ E.2 Phase 2: 5-tuple extraction
Figure F.1: Two DTKGs constructed using ATOM and Graphiti from 09-01-2020 (in UNIX, 1578524400) to 23-01-2020 (in UNIX, 1579734000) from 2020-COVID-NYT dataset. Left (ATOM): Preserves observation times ($t_{obs}$) separately from validity periods, with timestamps encoded in UNIX format to eliminate overhead associated with string parsing operations and timezone conversion calculations. Right (Graphiti): Treats observation time as validity start time. valid_atvalid\_at corresponds to $t_{\text{start}}$ in Graphiti’s time modeling. The highlighted fact “The mysterious respiratory virus spread to at least 10 other countries” is observed on 23-01-2020, but this does not guarantee the spread occurred at that time. ATOM’s dual-time modeling prevents such temporal misattribution.
@@
Figure F.2: Temporal resolution comparison between ATOM and Graphiti. Two atomic facts observed on January 28, 2020, report death counts from January 24 (26 deaths) and January 27 (at least 80 deaths). Left (ATOM): performs temporal resolution by detecting similar relations and extending their validity period history ($t_{\text{end}}$ in the figure). Right (Graphiti): creates separate relations for each atomic fact, resulting in duplication. Moreover, Graphiti misinterprets “By January 24, 2020” and “By January 27, 2020” as validity start times rather than validity end times, leading to temporal misattribution.
@@
Figure F.3: Temporal resolution comparison between ATOM and Graphiti. Two atomic facts observed on different dates (April 16 and April 19, 2020) describe protest activities during two time periods (“the week of April 13” and “the week of April 19”). Left (ATOM): merges similar relations and extend their validity periods ($t_{\text{start}}$ and $t_{\text{end}}$ in the figure). Right (Graphiti): maintains separate relations for each atomic fact. Moreover, Graphiti fails to translate “In the week of April 13, 2020” and “In the week of April 19, 2020” into proper validity periods as ATOM does.
Table T.2: Cost and token usage analysis on the 2020-COVID-NYT dataset. (L) denotes direct extraction from Lead paragraphs. (F) denotes the full ATOM pipeline, broken down into Module-1 (atomic fact decomposition) and Module-2 (5-tuple extraction from atomic facts).
Direct Extraction ATOM (F)
Metric (L) Module-1 Module-2 Total
Token Usage Statistics
Total Input Tokens 271,108 237,192 285,618 522,810
Total Output Tokens 83,757 38,646 217,110 255,756
Avg Input per Article 2,222 1,944 2,341 4,285
Avg Output per Article 687 317 1,780 2,096
Estimated Cost (USD)
claude-sonnet-4-2025-01-31 \$1.04 \$0.64 \$2.06 \$2.70
gpt-4.1-2025-04-14 \$0.61 \$0.40 \$1.15 \$1.55